A 0.2kg is drag 0.5m upward with a uniform velocity along a plane inclined 30° with the horizontal by a force parallel to the incline. The coefficient of kinetic friction between the object and the plane is 0.25. How much work is done by the applied force? Show complete solution.


1
Expert's answer
2021-12-02T10:04:15-0500

Let's first find the applied force :


"F_{appl}-F_{fr}-mgsin\theta=0,""F_{appl}=mgsin\theta+\mu_k mgcos\theta=mg(sin\theta+\mu_k cos\theta)."

Finally, we can find the work done by the applied force:


"W_{appl}=F_{appl}d=mgd(sin\theta+\mu_k cos\theta),""W_{appl}=0.2\ kg\times9.8\ \dfrac{m}{s^2}\times0.5\ m\times(sin30^{\circ}+0.25\times cos30^{\circ}),""W_{appl}=0.7\ J."